(2 points) Calculate the [OH-] of the following solutions: 2.6 x 10-5M HCl 0.20 M HNO3...

  1. (2 points) Calculate the [OH-] of the following solutions:
  1. 2.6 x 10-5M HCl
  2. 0.20 M HNO3
  3. 2.7 M x 10-9M HClO4
  4. 1.9 M HClO4
Answer #1

a)
use:
[OH-] = Kw/[H+]
Kw is dissociation constant of water whose value is 1.0*10^-14 at 25 oC
[OH-] = (1.0*10^-14)/[H+]
[OH-] = (1.0*10^-14)/(2.6*10^-5)
[OH-] = 3.846*10^-10 M
Answer: 3.8*10^-10 M

b)
use:
[OH-] = Kw/[H+]
Kw is dissociation constant of water whose value is 1.0*10^-14 at 25 oC
[OH-] = (1.0*10^-14)/[H+]
[OH-] = (1.0*10^-14)/(0.2)
[OH-] = 5.0*10^-14 M
Answer: 5.0*10^-14 M

c)
use:
[OH-] = Kw/[H+]
Kw is dissociation constant of water whose value is 1.0*10^-14 at 25 oC
[OH-] = (1.0*10^-14)/[H+]
[OH-] = (1.0*10^-14)/(2.7*10^-9)
[OH-] = 3.704*10^-6 M
Answer: 3.7*10^-6 M

d)
use:
[OH-] = Kw/[H+]
Kw is dissociation constant of water whose value is 1.0*10^-14 at 25 oC
[OH-] = (1.0*10^-14)/[H+]
[OH-] = (1.0*10^-14)/(1.9)
[OH-] = 5.263*10^-15 M
Answer: 5.3*10^-15 M